# Hölder curves with exotic tangent spaces

### v2

@@ 10.1. Self-similar sponges
First, if $Q_{k}\neq\emptyset$, then by Remark 10.1 $\operatorname{dim_{H}}(Q)=s$. Let $z\in Q_{k}$, let $T\in\operatorname{Tan}(Q,z)$ be bi-Lipschitz homeomorphic to $T^{*}_{k}$ for some $k\in\mathbb{N}$, and let $g:T^{*}_{k}\to T$ be a bi-Lipschitz homeomorphism. Then by Remark 8.12, there exists a countable collection of similarities $\{\phi_{j}:\mathbb{R}^{2}\to\mathbb{R}^{2}\}_{j\in\mathbb{N}}$ so that $T_{n,k}=\bigcup_{j\in\mathbb{N}}\phi_{j}(K^{n,0})$ and so that for every pair of distinct $i,j\in\mathbb{N}$, $\phi_{i}((0,1)^{2})\cap\phi_{j}((0,1)^{2})=\emptyset$. Then for each $j\in\mathbb{N}$, the composition $(g\circ I^{*}_{k}\circ\phi_{j}):K^{n,0}\to T$ is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism onto its image, further satisfying $(g\circ I^{*}_{k}\circ\phi_{j})(K^{n,0}\cap(0,1)^{2})$ is an open subset of $T$. Additionally, by Lemma 1.3 and its proof, for every $R>0$ there is an integer $M_{R}\in\mathbb{N}$ and a collection of similarities $f_{1},\dots,f_{M_{R}}:\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d^{\prime}}$ such that
$\overline{B}({\bf 0},R)\cap T\subset\bigcup_{j=1}^{M_{R}}f_{j}(Q)\subset\overline{B}({\bf 0},2R)\cap T$
and so that for every pair of distinct $i,j\in\{1,\dots,M_{R}\}$, $f_{i}(U)\cap f_{j}(U)=\emptyset$ and $f_{i}(U\cap Q)$ is an open subset of $T$.
Now let $i\in\mathbb{N}$ and $j\in\{1,\dots,M_{R}\}$ so that
@@
is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism. Recall the similarities $\psi^{2}_{n,1},\dots\psi_{n,5n-6}^{2}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from Section 4, the similarities generating the self-similar sponge $K^{n,0}$. Recall that $K^{n,0}\cap(0,1)^{2}\neq\emptyset$ and recall from the definition of self-similar sponge that the system $\{\psi_{n,1}^{2},\dots,\psi_{n,5n-6}^{2}\}$ satisfies the OSC via the open set $(0,1)^{2}$. Then there is a word $u\in\mathcal{A}_{n}^{*}$ so that
@@
All that remains is to prove the claim, and to this end let $y\in B$ and let $S\in\operatorname{Tan}(Q,y)$. Since $B$ is an open subset of $Q$, $S\in\operatorname{Tan}(B,y)$, so there exists a word $w\in\{1,\dots,m\}^{*}$ with $S\in\operatorname{Tan}(S_{w}(D),y)$. Note that the composition
@@
is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism, so by Lemma 9.1 the tangents $S$ and $\tilde{S}$ are $(\alpha_{n}/s)$-bi-Hölder homeomorphic. ∎
@@ Appendix A A Lipschitz curve and a point with extreme tangent space
Recall that $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$ denotes the collection of all sets $X\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ (i.e. closed sets in $\mathbb{R}^{N}$ that contain the origin $\bf 0$}) such that every component of $X$ is unbounded. In Lemma 2.5 we showed that for every non-constant Lipschitz curve in $\mathbb{R}^{N}$, the tangent space at every point is a subset of $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$. In this appendix we prove that the tangent space can in fact be all of $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$.
@@
We claim that $Y_{q}$ is connected. To this end, let $z_{0}\in Z_{q}$ be a point in $Z_{q}$ with minimal distance to $q$, and let $z\in Z_{q}$ be any other vertex. We now construct a chain of points $z_{0},z_{1},\dots,z_{M}\in Z_{q}$ such that $z_{M}=z$, and so that for every $i\in\{0,1,\dots,M-1\}$, $z_{i},z_{i+1}]\subset Y_{q}$. We proceed by induction, noting that the base vertex $z_{0}$ has already been established. Fix some integer $i\geq 0$ and assume we have already some point $z_{i}\in Z_{q}$, furthermore satisfying that if $i\geq 1$ then $z_{i-1},z_{i}]\subset Y_{q}$. Let $z_{i}=(z_{i}^{1},\dots,z_{i}^{N})$. If $z_{i}=z$, then we terminate the process. If not, then let $z=(z^{1},\dots,z^{N})$ and let $m\in\{1,\dots,N\}$ be the first index such that $z_{i}^{m}\neq z^{m}$. If $z_{i}^{m}<z^{m}$, then let $z_{i+1}^{m}=z_{i}^{m}+2^{-j}$, and if instead $z_{i}^{m}>z^{m}$, then let $z_{i+1}^{m}=z_{i}^{m}-2^{-j}$. For each $\ell\in\{1,\dots,N\}$ other than $\ell=m$, let $z_{i+1}^{\ell}=z_{i}^{\ell}$, and let $z_{i+1}=(z_{i+1}^{1},\dots,z_{i+1}^{N})$. Note that since $|z_{0}-q|$ is minimal, we have that for every $\ell\in\{1,\dots,N\}$, the difference in coordinates $|z_{0}^{\ell}-q^{\ell}|$ is also minimal among points in $Z_{q}$. Furthermore, by the construction of $z_{i+1}$, we have that for every $\ell\in\{1,\dots,N\}$, $z_{i+1}^{\ell}$ is between $z_{0}^{\ell}$ and $z^{\ell}$, possibly equal to one of these. Then $|z_{i}-z_{i+1}|=2^{-j}$ and $|z_{i+1}-q|\leq|z-q|\leq 2^{2-j}\sqrt{N}$, so $z_{i+1}\in Z_{q}$ and $z_{i},z_{i+1}]\subset Y_{q}$. This process terminates after finitely many steps, completing the construction, and so we have that $Y_{q}$ is connected.
Note that for points $q,u\in K$ which have $|q-u|\leq 2^{-j}$, we have that $Y_{q}\cap Y_{u}\neq\emptyset$. Now let $q\in K\cap\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}$, and since $K$ is connected there exists finite chain of points $\{p_{1},\dots,p_{M}\}\subset K$ so that $p_{1}=p$, $p_{M}=q$, and for every $i\in\{1,\dots,M-1\}$, $|p_{i}-p_{i+1}|\leq 2^{-j}$. Then $\bigcup_{i=1}^{M}Y_{p_{i}}\subset X_{j}$ is a connected set containing $v$ and intersecting with $\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}$, so it is contained in $B$, and therefore $B\cap\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}\neq\emptyset$. ∎
@@
then $(V^{j},E^{j})$ is connected by Lemma A.3, so it is equal to $G_{n_{j}}\in\mathcal{G}_{2j}$ for some $n_{j}\in\mathbb{N}$. Furthermore, $\text{Im}(G_{n_{j}})\cap[-1+4^{-j},1-4^{-j}]^{N}=2^{-j}X_{j}$, as desired. ∎
@@
Recall that $k_{n_{j}}=2j$ for each $j\in\mathbb{N}$, so $r_{n_{j}+1}\rho_{j}\leq 2^{-n_{j}-j+1}$ and $\rho_{j}r_{n_{j}}=2^{j}$. Hence, for every $j\in\mathbb{N}$,
@@
As $j\to\infty$, the first term of the latter sum approaches 0 by (A.1) and by design of the sets $G_{n_{j}}$. In the second term of the latter sum, the sets differ only inside $\rho_{j}r_{n_{j}+1}[-1,1]^{N}$, so the excess is at most $\operatorname{diam}(\rho_{j}r_{n_{j}+1}[-1,1]^{N})\leq 2\rho_{j}r_{n_{j}+1}\sqrt{N}$, which approaches 0 as $j\to\infty$. This completes the proof, by definition of convergence $\lim_{j\to\infty}\rho_{j}H$ in the Attouch-Wets topology. ∎

### v3

@@ 10.1. Tangents of self-similar sponges
First, if $Q_{k}\neq\emptyset$, then by Proposition 10.1, $\operatorname{dim_{H}}(Q)=s$. Let $z\in Q_{k}$, let $T\in\operatorname{Tan}(Q,z)$ be bi-Lipschitz homeomorphic to $T^{*}_{k}$ for some $k\in\mathbb{N}$, and let $g:T^{*}_{k}\to T$ be a bi-Lipschitz homeomorphism. Then by Remark 8.12, there exists a countable collection of similarities $\{\phi_{j}:\mathbb{R}^{2}\to\mathbb{R}^{2}\}_{j\in\mathbb{N}}$ such that $T_{n,k}=\bigcup_{j\in\mathbb{N}}\phi_{j}(K^{n,0})$ and such that for every pair of distinct $i,j\in\mathbb{N}$, $\phi_{i}((0,1)^{2})\cap\phi_{j}((0,1)^{2})=\emptyset$. Then for each $j\in\mathbb{N}$, the composition $(g\circ I^{*}_{k}\circ\phi_{j}):K^{n,0}\to T$ is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism onto its image, further satisfying $(g\circ I^{*}_{k}\circ\phi_{j})(K^{n,0}\cap(0,1)^{2})$ is an open subset of $T$. Additionally, by Lemma 10.1 and its proof, for every $R>0$ there is an integer $M_{R}\in\mathbb{N}$ and a collection of similarities $f_{1},\dots,f_{M_{R}}:\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d^{\prime}}$ such that
$\overline{B}({\bf 0},R)\cap T\subset\bigcup_{j=1}^{M_{R}}f_{j}(Q)\subset\overline{B}({\bf 0},2R)\cap T,$
further satisfying for every pair of distinct $i,j\in\{1,\dots,M_{R}\}$, $f_{i}(U)\cap f_{j}(U)=\emptyset$ and $f_{i}(U\cap Q)$ is an open subset of $T$.
Now let $i\in\mathbb{N}$ and $j\in\{1,\dots,M_{R}\}$ such that
@@
is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism. Recall the similarities $\psi^{2}_{n,1},\dots\psi_{n,5n-6}^{2}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from Section 4, the similarities generating the self-similar sponge $K^{n,0}$. Recall that $K^{n,0}\cap(0,1)^{2}\neq\emptyset$ and recall from the definition of a self-similar sponge that the system $\{\psi_{n,1}^{2},\dots,\psi_{n,5n-6}^{2}\}$ satisfies the OSC via the open set $(0,1)^{2}$. Then there is a word $u\in\mathcal{A}_{n}^{*}$ satisfying
@@
All that remains is to prove the claim, and to this end let $y\in B$ and let $S\in\operatorname{Tan}(Q,y)$. Since $B$ is an open subset of $Q$, $S\in\operatorname{Tan}(B,y)$, thus there exists a word $w\in\{1,\dots,m\}^{*}$ with $S\in\operatorname{Tan}(S_{w}(D),y)$. Note that the composition
@@
is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism, therefore by Lemma 9.1 the tangents $S$ and $\tilde{S}$ are $(\alpha_{n}/s)$-bi-Hölder homeomorphic. ∎
@@ Appendix A A Lipschitz curve and a point with extreme tangent space
Let $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$ denote the collection of all closed sets $X\subset\mathbb{R}^{N}$ that contain the origin $\bf 0$} such that every component of $X$ is unbounded. In Lemma 2.5 we showed that for every non-constant Lipschitz curve in $\mathbb{R}^{N}$, the tangent space at every point is a subset of $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$. In this appendix we prove that the tangent space can in fact be all of $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$.
@@
We claim that $Y_{q}$ is connected. To this end, let $z_{0}\in Z_{q}$ be a point in $Z_{q}$ with minimal distance to $q$, and let $z\in Z_{q}$ be any other vertex. We now construct a chain of points $z_{0},z_{1},\dots,z_{M}\in Z_{q}$ such that $z_{M}=z$, and such that for every $i\in\{0,1,\dots,M-1\}$, $z_{i},z_{i+1}]\subset Y_{q}$. We proceed by induction, noting that the base vertex $z_{0}$ has already been established. Fix some integer $i\geq 0$ and assume we have already some point $z_{i}\in Z_{q}$, furthermore satisfying that if $i\geq 1$ then $z_{i-1},z_{i}]\subset Y_{q}$. Let $z_{i}=(z_{i}^{1},\dots,z_{i}^{N})$. If $z_{i}=z$, then we terminate the process. If not, then let $z=(z^{1},\dots,z^{N})$ and let $m\in\{1,\dots,N\}$ be the first index such that $z_{i}^{m}\neq z^{m}$. If $z_{i}^{m}<z^{m}$, then let $z_{i+1}^{m}=z_{i}^{m}+2^{-j}$, and if instead $z_{i}^{m}>z^{m}$, then let $z_{i+1}^{m}=z_{i}^{m}-2^{-j}$. For each $\ell\in\{1,\dots,N\}$ other than $\ell=m$, let $z_{i+1}^{\ell}=z_{i}^{\ell}$, and let $z_{i+1}=(z_{i+1}^{1},\dots,z_{i+1}^{N})$. Note that since $|z_{0}-q|$ is minimal, we have that for every $\ell\in\{1,\dots,N\}$, the difference in coordinates $|z_{0}^{\ell}-q^{\ell}|$ is also minimal among points in $Z_{q}$. Furthermore, by the construction of $z_{i+1}$, we have that for every $\ell\in\{1,\dots,N\}$, $z_{i+1}^{\ell}$ is between $z_{0}^{\ell}$ and $z^{\ell}$, possibly equal to one of these. Then $|z_{i}-z_{i+1}|=2^{-j}$ and $|z_{i+1}-q|\leq|z-q|\leq 2^{2-j}\sqrt{N}$, hence $z_{i+1}\in Z_{q}$ and $z_{i},z_{i+1}]\subset Y_{q}$. This process terminates after finitely many steps, completing the construction, and therefore $Y_{q}$ is connected.
Note that for points $q,u\in K$ which have $|q-u|\leq 2^{-j}$, we have that $Y_{q}\cap Y_{u}\neq\emptyset$. Now let $q\in K\cap\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}$, and since $K$ is connected there exists finite chain of points $\{p_{1},\dots,p_{M}\}\subset K$ such that $p_{1}=p$, $p_{M}=q$, and for every $i\in\{1,\dots,M-1\}$, $|p_{i}-p_{i+1}|\leq 2^{-j}$. Then $\bigcup_{i=1}^{M}Y_{p_{i}}\subset X_{j}$ is a connected set containing $v$ and intersecting with $\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}$, thus it is contained in $B$, and therefore $B\cap\partial[2^{-j}-2^{j},2^{j}-2^{-j}]^{N}\neq\emptyset$. ∎
@@
then $(V^{j},E^{j})$ is connected by Lemma A.3, hence it is equal to $G_{n_{j}}\in\mathcal{G}_{2j}$ for some $n_{j}\in\mathbb{N}$. Furthermore, $\text{Im}(G_{n_{j}})\cap[-1+4^{-j},1-4^{-j}]^{N}=2^{-j}X_{j}$, as desired. ∎
@@
Recall that $k_{n_{j}}=2j$ for each $j\in\mathbb{N}$, thus $r_{n_{j}+1}\rho_{j}\leq 2^{-n_{j}-j+1}$ and $\rho_{j}r_{n_{j}}=2^{j}$. Hence, for every $j\in\mathbb{N}$,
@@
As $j\to\infty$, the first term of the latter sum approaches 0 by (A.1) and by design of the sets $G_{n_{j}}$. In the second term of the latter sum, the sets differ only inside $\rho_{j}r_{n_{j}+1}[-1,1]^{N}$, therefore the excess is at most $\operatorname{diam}(\rho_{j}r_{n_{j}+1}[-1,1]^{N})\leq 2\rho_{j}r_{n_{j}+1}\sqrt{N}$, which approaches 0 as $j\to\infty$. This completes the proof, by definition of convergence $\lim_{j\to\infty}\rho_{j}H$ in the Attouch-Wets topology. ∎